# FMANet: A Novel Dual-Phase Optical Flow Approach with Fusion Motion Attention Network for Robust Micro-expression Recognition

### v1

@@ I Introduction
Input modality in MER can be broadly categorized into appearance-based and motion-based approaches. Appearance-based methods rely on static features extracted from individual frames, such as texture, shape, and facial landmarks. While these methods can capture spatial information, they often struggle to effectively detect the fleeting and low-intensity movements that define micro-expressions. In contrast, motion-based approaches analyze facial movement patterns directly, making them more effective in recognizing micro-expressions [6]. Among motion-based methods, optical flow-based techniques have gained prominence due to its ability to capture subtle facial movements while minimizing identity-specific features [7, 8]. Various studies have explored different ways to leverage optical flow for micro-expression analysis, each with its own strengths and limitations. Early approaches such as Bi-WOOF [9] and MDMO [10] primarily relied on handcrafted features extracted from optical flow. Bi-WOOF (Binarized Weighted Oriented Optical Flow) encoded directional movement information to enhance discriminability, while MDMO (Main Directional Mean Optical Flow) focused on aggregating dominant motion directions. Despite their effectiveness in highlighting key motion patterns, these methods suffered from limited representational capacity and were highly sensitive to noise in low-texture regions. Unlike the aforementioned works which exploited only the single dominant direction of optical flow in each facial region, Allaert et al. [11] introduced a refined motion representation by emphasizing coherent optical flow structures. However, these approaches often involved complex architectures, increasing computational costs and requiring large-scale datasets for effective training. Liong et al. [12] proposed OFF-ApexNet, a learning-based approach that extracts optical flow features specifically at the apex frame to reduce redundant temporal information.
@@ II-A Optical Flow Calculation
To estimate the optical flow, it is generally assumed that:
  • The brightness of the moving objects remains unchanged between the source and target frames.
  • The displacement of each pixel between consecutive frames is minimal, as object motion occurs gradually over time.
  • The background and objects belonging to it are assumed to be static.
Micro-expression (ME) characteristics align well with these assumptions, making optical flow a suitable choice for ME analysis:
  • ME datasets are collected under strictly controlled lighting conditions.
  • The motion intensity in MEs is extremely subtle, often imperceptible to the naked eye, and occurs over a very short duration $(<0.5s)$, ensuring minimal displacement between consecutive frames.
  • The data collection process ensures a fixed and stable background.
These properties make optical flow highly effective for MER, outperforming other motion-based methods.
@@ III-2 Soft Motion Attention
The second module, Soft Motion Attention Block (SMAB), reformulates magnitude modulation into a learnable, soft attention mechanism. It refines the fused feature map from the FFB by selectively amplifying salient motion regions and suppressing noise. Instead of applying thresholds $(\alpha,\beta)$ and weights $(w_{1},w_{2})$, The attention mechanism in SMAB is guided by two motion-derived cues: the joint magnitude (capturing overall motion strength - $M_{c}$) and the directional coherence (measuring consistency between phases - $C$). The combined magnitude captures the joint motion intensity of both phases using a geometric mean:
@@ IV-A Dataset
  • CASME-II: Developed by the Chinese Academy of Sciences, the dataset consists of a total of 255 videos collected from 26 participants. It categorizes videos into seven emotion groups: others (99 videos), disgust (63 videos), happiness (32 videos), repression (27 videos), surprise (25 videos), sadness (7 videos), and fear (2 videos).
  • SAMM: It includes 159 spontaneous expression videos collected from 32 participants. The dataset categorizes expressions into 8 types as follows: anger (57 videos), happiness (26 videos), other (26 videos), surprise (15 videos), contempt (12 videos), disgust (9 videos), fear (8 videos) and sadness (6 videos).
  • MMEW: The MMEW dataset includes 300 spontaneous micro-expression samples collected in more naturalistic settings. Videos were recorded at 90 fps and annotated with senven emotion categories: happiness (36 videos), anger (8 videos), surprise (89 videos), disgust (72 videos), fear (16 videos), sadness (13 videos), others (66 videos). MMEW introduces more variation in head pose and lighting, making it a challenging benchmark for generalization.
@@ IV-E3 Analysis of Magnitude-modulated Combined Optical Flow (MM-COF)
Input modalities: Finally, Table XI compares MM-COF with other representations. It is observed that magnitude-modulated representations (MM-OF and MM-COF) consistently outperform plain optical flow and apex frame inputs. Among the backbones, SCNN delivers higher and more stable performance than VGG19 across both datasets, confirming that a compact, task-specific architecture is better suited for capturing subtle micro-expression motions than a generic deep model. Notably, MM-COF with SCNN achieves the best single-stream results, reaching 70.08% on CASME-II and 63.70% on SAMM.

### v2

@@ I Introduction
Input modality in MER can be broadly categorized into appearance-based and motion-based approaches. Appearance-based methods rely on static features extracted from individual frames, such as texture, shape, and facial landmarks. While these methods can capture spatial information, they often struggle to effectively detect the fleeting and low-intensity movements that define micro-expressions. In contrast, motion-based approaches analyze facial movement patterns directly, making them more effective in recognizing micro-expressions [6]. Among motion-based methods, optical flow-based techniques have gained prominence due to its ability to capture subtle facial movements while minimizing identity-specific features [7, 8]. Various studies have explored different ways to leverage optical flow for micro-expression analysis, each with its own strengths and limitations. Early approaches such as Bi-WOOF [9] and MDMO [10] primarily relied on handcrafted features extracted from optical flow. Despite their effectiveness in highlighting key motion patterns, these methods suffered from limited representational capacity and were highly sensitive to noise in low-texture regions. Unlike the aforementioned works which exploited only the single dominant direction of optical flow in each facial region, Allaert et al. [11] introduced a refined motion representation by emphasizing coherent optical flow structures. However, these approaches often involved complex architectures, increasing computational costs and requiring large-scale datasets for effective training.
@@ III-2 Soft Motion Attention
The second module, Soft Motion Attention Block, reformulates magnitude modulation into a learnable, soft attention mechanism. It refines the fused feature map from the FFB by selectively amplifying salient motion regions and suppressing noise. Instead of applying thresholds $(\alpha,\beta)$ and weights $(w_{1},w_{2})$, The attention mechanism in SMAB is guided by two motion-derived cues: the joint magnitude (capturing overall motion strength - $M_{c}$) and the directional coherence (measuring consistency between phases - $C$). The combined magnitude captures the joint motion intensity of both phases using a geometric mean:
@@ IV-C Evaluation Metrics
Figure 6: Onset frame, Apex frame, Offset frame, Single Optical Flow, Combined Optical Flow and Magnitude-Modulated Combined Optical Flow (ordered from left to right) on CASME-II and SAMM datasets. (Best viewed in color)
@@ IV-D Results
IV-D1 Visual representation results
Fig. 6 demonstrate the differences between Single Optical Flow, Combined Optical Flow, and our proposed Magnitude-Modulated Combined Optical Flow, along with the onset, apex, and offset frames. By applying the threshold, our method focuses on the regions with significant motion, filtering out irrelevant or less dynamic areas. Visual inspection shows that the MM-COF captures more detailed transitions in facial expressions, particularly those that are crucial for recognizing micro-expressions. This selective focus on key dynamic regions likely contributes to the observed performance improvement in Micro-Expression Recognition compared to Single Optical Flow and Combined Optical Flow.
Figure 7: Optical flow and the corresponding FFB Features map and SMAB Attention map. (Best viewed in color)
Figure 7 visualizes the outputs of the components within the proposed FMANet architecture. From left to right, the images represent: the optical flow during the onset phase, the optical flow during the offset phase, the corresponding feature map generated by the FFB (Phase-Aware Consensus Fusion Block), and the attention map obtained from the SMAB (Soft Motion Attention Block).
The results indicate that both the FFB and SMAB blocks successfully fulfill their intended purpose. The FFB effectively combines homogeneous regions from the two motion phases, and the SMAB concentrates on the critical regions within the FFB’s feature representation. Specifically, compared to the original optical flow representations, the motion features generated by the FFB are more effective at reducing noise and focusing on significant areas. For example, in the ”Surprise” and ”Happiness” expressions, the movements are subtle and tend to produce considerable noise in the optical flow representation. However, our method successfully distills the core motion information by unifying the two motion phases, and the SMAB effectively targets these key regions.
Furthermore, the results demonstrate that with the representations generated by the FFB and SMAB, the corresponding Action Units can be observed and identified much more easily and accurately than with the original optical flow representation.
IV-D2 Comparison with State-of-the-arts
@@ IV-E3 Analysis of Magnitude-modulated Combined Optical Flow (MM-COF)
Input modalities: Finally, Table XI compares MM-COF with other representations. It is observed that magnitude-modulated representations (MM-COF) consistently outperform plain optical flow and apex frame inputs. Among the backbones, SCNN delivers higher and more stable performance than VGG19 across both datasets, confirming that a compact, task-specific architecture is better suited for capturing subtle micro-expression motions than a generic deep model. Notably, MM-COF with SCNN achieves the best single-stream results, reaching 70.08% on CASME-II and 63.70% on SAMM.